OpenStudy (vianne):

What is the estimated value of the expression if each number in the expression is rounded to the nearest whole unit? (6.82)(2.07)(4.35)

OpenStudy (vianne):

@waterineyes

OpenStudy (anonymous):

What do you say?

OpenStudy (anonymous):

Any try you gave to it?

OpenStudy (anonymous):

\(6.82\) is between \(6\) and \(7\), and tell me it is nearer to 6 or 7?

OpenStudy (vianne):

I did try.. I just don't know what I'm looking for. :/

OpenStudy (vianne):

7

OpenStudy (anonymous):

And 2.07 ?

OpenStudy (anonymous):

nearer to \(2\) or \(3\)?

OpenStudy (vianne):

Closer to 2.1

OpenStudy (vianne):

2

OpenStudy (vianne):

4.35 is closer to 4.4/4

OpenStudy (vianne):

So the answer would be (7)?

OpenStudy (anonymous):

And same do for last one.

OpenStudy (vianne):

(7)(2)(4)?

OpenStudy (anonymous):

so, you are done.

OpenStudy (anonymous):

As simple as that.. :)

OpenStudy (vianne):

I'm literally stupid, do I type 7 with parentheses or just 7 solo? ._. There's no answer choices just a box to fill in :3

OpenStudy (anonymous):

May I see the screenshot of that where you are going to write this?

OpenStudy (anonymous):

need not to use parenthesis, you can use multiply sign in between, but if multiply sign is not allowed then you must use brackets. :)

OpenStudy (anonymous):

Wait.. What is the final result you got?

OpenStudy (anonymous):

(7)(2)(4) is not the final result.. what you can do more in this?

OpenStudy (vianne):

7 x 2 x 4?

OpenStudy (vianne):

Actually round 7? :3

OpenStudy (anonymous):

What else? you can do more..

OpenStudy (anonymous):

What is \(7 \times 2 = ??\)

OpenStudy (vianne):

14

OpenStudy (anonymous):

Then , \(7 \times 2 \times 4 = ??\)

OpenStudy (vianne):

56

OpenStudy (anonymous):

This is your final result, this is the question to find out the final estimated value of given expression.. :)

OpenStudy (vianne):

ohh I was so confused on that part, since I rounded everything, just.. Didn't know what to put into the answer box.

OpenStudy (anonymous):

Now you know?

OpenStudy (vianne):

Yup

OpenStudy (anonymous):

Observe, \(56 = \text{your Smart Score - 1}\) :P

OpenStudy (anonymous):

\(\text{This is it...}\)

OpenStudy (vianne):

:P

OpenStudy (vianne):

This will be good use for a future 7th grader trying to solve this question out. :3

OpenStudy (vianne):

100% since I did the other 4 myself.

OpenStudy (anonymous):

Excellent.. Going great.. :)